OpenStudy (anonymous):

How many solutions (using only nonnegative integers) are there to the following equation? x1 + x2 + x3 + x4 + x5 = 35

OpenStudy (loser66):

hello, sweeties.

OpenStudy (loser66):

we have formula for this problem, look it up at your note, it's there.

OpenStudy (loser66):

C(n +r -1, r) where n is number of elements, here n =5, r =35 so, you have C( 5+35-1, 35) = C( 39,35)= C( 39,4) \[= \frac{39*38*37*36}{4*3*2*1}=82251\]

OpenStudy (anonymous):

thx love